# Dynamical Phase Transitions, Caustics, and Quantum Dark Bands

### v1

@@ Abstract
We provide a new perspective on dynamical phase transitions (DPTs) in quantum systems by showing their connection to caustics forming in the Fock space representation of the many-body state over time. In this way we find an explanation for DPTs in a quantum spin system in terms of an everyday natural phenomenon: The dark band between the primary and seconday bows (caustics) in rainbows (Alexander’s dark band). The DPT occurs when the Loschmidt echo crosses the switching line between the evanescent tails of two back-to-back Airy functions that dress neighbouring fold caustics in Fock space and is the time-dependent analogue of what is seen as a function of angle in the sky. The structural stability and universal properties of caustics, as described mathematically by catastrophe theory, suggests our analysis has wide applicability. Based on our thorough analytical understanding we propose a protocol which can be used to verify the existence of a DPT in a finite system experiment.
@@ I Introduction
Phase transitions are characterised by singular changes in the properties of a macroscopic system as a control parameter is varied smoothly. For example, the well-known mathematical treatment of phase transitions by Yang and Lee shows how the free energy in the Ising model develops a logarithmic singularity at a critial value of the temperature as the thermodynamic limit is approached [1, 2]. This non-analyticity is the ultimate source of the self-similar scaling laws and universality that occurs across very different physical systems close to continuous phase transitions. The traditional paradigm of phase transitions concerns changes to the equilibrium state, which in the case of quantum phase transitions usually implies the ground state [3]. However, the study of more general non-equilibrium phase transitions that occur in driven-dissipative quantum systems, for instance in quantum optics, also has a long history [4, 5].
More recent developments are connected to the search for universality in general quantum dynamics [6, 7, 8, 9, 10]. One variety is known as excited state quantum phase transitions (ESQPTs) [11, 12, 13, 14, 15, 16] that arise from features in the quantum energy level distribution, such as the diverging density of states that occurs in the semiclassical regime at energies corresponding to a separatrix in classical phase space. A physical example of this is provided by the transition to macroscopic quantum self trapping in coupled atomic Bose-Einstein condensates where time-oscillating population differences become locked above a critical magnitude [17, 18, 19, 20, 21, 22, 23, 24]. The control parameter in ESQPTs is the initial energy relative to the ground state and is typically set by a well controlled initial quench in which a parameter in the Hamiltonian is suddenly changed, thereby generating non-equilibrium dynamics.
@@
In this paper we focus on another type of non-equilibrium phase transition where the control parameter is time itself. In this case, which has been categorized by Corps et al [25] as a type II dynamical phase transition, an observable changes non-analytically at a finite critical time following the initial sudden quench. We shall refer to this scenario simply as a dynamical phase transition (DPT). A DPT is clearly different to an ESQPT where, once launched at a particular energy, the system remains in the same phase over time (although there can be situations where there is an initial transient before the phase becomes well established, this is distinct from a non-analytic transition [16, 25]).
@@
DPTs have been studied mostly in the context of quantum many-body systems [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], although they also appear in classical physics [42, 34, 43]. In contrast to equilibrium transitions, which give rise to a global reorganization of the system that can be detected by macroscopic observables such as the magnetization of a bulk metal, the DPT phenomenon is rather subtle and difficult to detect directly in experiments because it involves measuring an exponentially suppressed quantity such as an overlap of many-body quantum states or the exponential tail of a classical probability distribution. In keeping with most works on DPTs, we focus in this paper on the probability to return to the initial state via the dynamics after time $t$. In the quantum case the return probability is known as the Loschmidt echo. Specifically, let $|\psi(t)\rangle$ be the time evolved state of a many-body system consisting of $N$ particles. Then the Loschmidt echo is given by
@@
The overlap of two distinct many-body states is generally expected to be exponentially small in the system size, a property known as the ‘orthogonality catastrophe’—almost all states in Hilbert space become orthogonal in the large-$N$ limit. This makes experimentally measuring the Loschmidt echo very difficult in a many-body context. Mathematically, we can describe the Loschmidt echo of a many-body system with the rate function $r(t)$ which typically has a finite limit. DPTs are characterized by nonanalytic behavior of this function.
In this article we focus on the fully-connected version of the well-known transverse-field Ising model (TFIM). The model will be introduced in detail later, but we emphasise here that it has been used as the minimal model for describing dynamics in a wide range of physical systems including bosonic Josephson junctions in ultracold atomic gases [19, 20, 21, 22, 23, 24], weak links in superfluid 3He [44], ferroelectrics such as KH2PO4 and ferromagnets such as Dy(C2H5SO4)39H2O [45], exciton polaritons in semiconductors [46, 47], and light beams carrying angular momentum [48]. The model is also relevant to TFIM systems that are not fully connected but have long range interactions such as those realized using trapped atomic ions [49, 50, 51, 30, 31], and the Dicke model describing cold atoms in optical cavities [52, 53, 54, 55]. A typical example of the dynamics of the rate function for our model is displayed in Fig. 1. We see that the return function develops a kink as the system size is increased, and the limiting function for $N\rightarrow\infty$ is clearly nonanalytic. This paradigmatic ‘sharkfin’ shape is common to many DPTs [28]. Such behavior cannot be imposed from the mean-field theory even though it accurately describes the evolution of macroscopic observables. Instead, the Loschmidt echo is determined by weak quantum fluctuations that are exponentially suppressed in fully-connected systems.
@@
Figure 1: Paradigmatic example of a dynamical phase transition. The Loschmidt rate function acquires a nonanalytic kink in the limit of large system size $N\rightarrow\infty$ at a finite time [56]. The model used here and throughout this paper is the fully-connected transverse field Ising model given in Eq. (4). In this example we have choosen the ground state at zero transverse field ($\Omega=0$) as the initial state, which we denote as $|\psi(0)\rangle=|0\rangle^{\otimes N}$, and at $t=0$ we quench the transverse field from $\Omega=0$ to $\Omega=G$.
DPTs have been studied in the fully connected transverse-field Ising model (FCTFIM) before [57, 56, 58, 59, 25]. Our main goal here is to point out the role played by caustics in Fock space and to make a connection to the phenomenon of dark bands between rainbows. Caustics are usually associated with waves or rays in real space and are the most intense regions of a wave-field. In geometric optics they are the envelopes of families of rays and form quite generically through natural focusing, e.g. optical rainbows focused by water droplets in the sky and gravitational lensing by mass distributions in the cosmos [60]. Similarly, freak waves, tidal bores and tsunamis are the result of the natural focusing of water waves by the topography of the sea bed or estuary shape [61, 62, 63, 64, 65]. Caustics are singularities of the classical theory: they are regions where the wave amplitude diverges in the short wavelength (ray) limit $\lambda\rightarrow 0$, and their natural mathematical description is in terms of catastrophe theory [60, 66]. Caustics in Fock space are the extension of these ideas to many-body quantum mechanics. A single ray in Fock space represents a single classical field configuration, while quantum configurations can be represented by the superposition of many rays. Whereas ordinary caustics are resolved by going to the wave theory where interference removes the singularity, caustics in Fock space are resolved by second-quantization where the discreteness of field excitations is the new feature needed to remove the singularity. The theory of Fock-space caustics has been developed over the past decade in the context of sudden quenches in cold atom systems [67, 68, 69, 70, 71] and is also relevant to dissipative systems [72].
Catastrophe theory makes the remarkable prediction that there are only a finite number of distinct shapes that structurally stable singularities can take, depending on the dimension of the space in which they live [73, 74]. Structural stability means that the singularity is qualitatively stable against perturbations and thus catastrophe theory deals with singularities that appear generically in nature because they do not require fine tuning to occur. In the two dimensions which are relevant to the dynamics of the FCTFIM, namely the one dimension of Fock space that gives the total magnetization (difference between the number of up and down spins) plus time, the only structurally stable catastrophes are fold lines and cusp points (fold lines meet at cusp points). Furthermore, in the wave theory each type of catastrophe is dressed by its own characteristic local wavefunction known as a ‘wave catastrophe’ or ‘diffraction integral’ [75, 76]. For the fold this is the Airy function and for the cusp it is the lesser-known Pearcey function. Both of these wavefunctions can be seen occurring in the left hand panel of Fig. 2 that shows the development of a many-body wavefunction in Fock space as a function of time. Higher dimensional models display higher wave catastrophes [71] that form a hierarchy with the higher ones containing the lower ones. Each type of catastrophe in fact forms an equivalence class with its own self-similar scaling properties and scaling indices that are the analogues of the universality classes and critical exponents of equilibrium phase transitions [66]. Catastrophe theory therefore reveals a type of universality in quantum dynamics due to caustics. We show in this work how dynamical phase transitions can be explained in this context.
@@
The plan for the rest of this paper is as follows. In Section II we specify the model and give an overview of our main results. This includes an explanation of Alexander’s dark band (DB) which is an everyday optical phenomenon visible in the sky between the primary and secondary bows of a rainbow [77] and turns out to be closely analogous to DPTs but with the role of time played by angle in the sky. Section III gives a more indepth treatment of rainbows, and Section IV presents the main theoretical calculation in the paper which gives the rate function in the semiclassical regime. Section V discusses a possible measurement scheme, which is a tricky issue because the Loschmidt echo is exponentially small at the DPT. Our solution is to perform two different measurements on two halves of the system. We give a general discussion of switching lines which are the essence of DPTs in Section VI, and our conclusions are set out in Section VII.
II Main results: Dynamical Phase Transitions and Caustics
@@
The model we use for studying dynamical phase transitions is the FCTFIM [56, 25]. It consists of $N$ two-level systems (Pauli operators $\vec{\sigma}^{i}$) with Hamiltonian
@@
These are commonly known as the Dicke states. However, we will refer to them as Fock states because they are the symmetrized single particle eigenstates of the number difference operator (giving the difference between the number of up and down spins, i.e. the magnetization). For the sake of working with a specific example, in what follows we set $\Omega=G$, and choose the initial state to be a single Fock state $|\psi_{0}\rangle=|j,m_{0}\rangle$, with $m_{0}$ chosen as the closest integer (or half integer) to $0.6j$. We emphasise that although a single Fock state might not appear to be a ‘typical’ initial condition, it is in fact not special because small variations in the initial state or the parameters in the Hamiltonian do not lead to qualitatively different outcomes, as is to be expected from the structural stability of catastrophes. For example, choosing a narrow gaussian wavepacket of Fock states still leads to similar fold and cusp patterns [69, 78]. In fact, going to the opposite situation of an initial phase state (which is very broad in Fock space) also gives rise to caustics [71], although of course they will be in different positions. The key ingredient for caustics to appear is that the initial state must be far above the ground state and such states can easily be generated in sudden quenches.
@@
Figure 2: Left: The modulus of the exact quantum wave function plotted in Fock space as a function of time. The shading gives the magnitude of the overlap $|\langle j,m|\psi(t)\rangle|$ with each Fock state. The parameters are $\Omega=G$, $j=80$, and the initial state is the single Fock state $|j=80,m=48\rangle$. The envelopes separating the bright and dark regions are fold caustics, and these meet at cusp caustics (in order to faciliate comparison with optical rainbows we use the terminology ‘bright’ and ‘dark’ although in the image colormap higher amplitudes corresponds to blacker shades and lower amplitudes to whiter shades). In the full quantum theory shown here, folds are dressed by Airy functions whose characteristic fringe pattern can be seen near the edges of the shaded (‘bright’) regions. Cusps are dressed by Pearcey functions as occurs near $m/j=-0.6$ and time $\Omega t\approx 2.4$. Right: The Loschmidt echo $L(t)$, or return probability, is given by the cut through the general wave function corresponding to the initial state. The cut is depicted by the horizontal blue line in the left panel. The Loschmidt echo shown here features a one dimensional slice through two fold caustics that bound a dark forbidden region where the wave function is exponentially small.

### v2

@@ Abstract
We provide a new perspective on quantum dynamical phase transitions (DPTs) by explaining their origin in terms of caustics that form in the Fock space representation of the many-body state over time, using the fully connected transverse field Ising model as an example. In this way we establish a connection between DPTs in a quantum spin system and an everyday natural phenomenon: The dark band between the primary and seconday bows (caustics) in rainbows known as Alexander’s dark band. The DPT occurs when the Loschmidt echo crosses the switching line between the evanescent tails of two back-to-back Airy functions that dress neighbouring fold caustics in Fock space and is the time-dependent analogue of what is seen as a function of angle in the sky. The structural stability and universal properties of caustics, as described mathematically by catastrophe theory, explains the generic occurrence of DPTs in the model and suggests that our analysis has wide applicability. Based on our thorough analytical understanding we propose a protocol which can be used to verify the existence of a DPT in a finite system experiment.
@@ I Introduction
Phase transitions are characterised by singular changes in the properties of a macroscopic system as a control parameter is varied smoothly [1, 2]. This non-analyticity is the source of the self-similar scaling laws and universality that occurs across very different physical systems close to continuous phase transitions. Whereas the traditional paradigm of phase transitions concerns changes to the equilibrium state, which in the case of quantum phase transitions usually implies the ground state [3], the study of more general non-equilibrium phase transitions that occur in driven-dissipative quantum systems also has a long history [4, 5]. More recent developments are connected to the search for universality in general quantum dynamics [6, 7, 8, 9, 10], and include excited state quantum phase transitions (ESQPTs) [11, 12, 13, 14, 15, 16] that arise from the diverging density of states that occurs in the semiclassical regime at energies corresponding to a separatrix in classical phase space [17, 18, 19, 20, 21, 22, 23, 24]. The control parameter in ESQPTs is the energy relative to the ground state and can be set by a well controlled initial quench in which a parameter in the Hamiltonian is suddenly changed.
@@
In this paper we focus on another type of non-equilibrium phase transition where the control parameter is time itself. In this case, which has been categorized by Corps et al [25] as a type II dynamical phase transition, an observable changes non-analytically at a finite critical time following the initial quench. We refer to this scenario simply as a dynamical phase transition (DPT). A DPT is different to an ESQPT where, once launched at a particular energy, the system remains in the same phase over time (although there can be situations where there is an initial transient [16, 25]).
@@
DPTs have been primarily studied in the context of quantum many-body systems, both theoretically [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] and experimentally [38, 39, 40], although they also appear in classical physics [41, 42, 43]. In contrast to equilibrium transitions, which give rise to a global reorganization of the system that can be detected by macroscopic observables such as the magnetization of a bulk metal, the DPT phenomenon is rather subtle and difficult to detect directly in experiments because it involves measuring an exponentially suppressed quantity such as an overlap of many-body quantum states or the exponential tail of a classical probability distribution. In keeping with most works on DPTs, we focus in this paper on the probability to return to the initial state via the dynamics after time $t$. In the quantum case the return probability is known as the Loschmidt echo. Specifically, let $|\psi(t)\rangle$ be the time evolved state of a many-body system consisting of $N$ particles. Then the Loschmidt echo is given by
@@
The overlap of two distinct many-body states is generally expected to be exponentially small in the system size, a property known as the ‘orthogonality catastrophe’—almost all states in Hilbert space become orthogonal in the large-$N$ limit. This makes experimentally measuring the Loschmidt echo very difficult in a many-body context. Mathematically, we can describe the Loschmidt echo of a many-body system with the rate function $r(t)$ (see Eq. (1)) which typically has a finite limit as $N\rightarrow\infty$. DPTs are characterized by nonanalytic behaviour of this function. A typical example of the dynamics of the rate function is displayed in Fig. 1. We see that the return function develops a kink as the system size is increased, and the limiting function for $N\rightarrow\infty$ is clearly nonanalytic. This characteristic ‘sharkfin’ shape is common to many DPTs, yet, the shape of the curve and the position of the kink (critical time) depend on the specific model in a highly non-trivial way [28].
@@
Figure 1: Paradigmatic example of a dynamical phase transition. The Loschmidt rate function acquires a nonanalytic kink in the limit of large system size $N\rightarrow\infty$ at a finite time [44]. The model used here and throughout this paper is the fully-connected transverse field Ising model given in Eq. (4). In this example we have choosen the ground state at zero transverse field ($\Omega=0$) as the initial state, which we denote as $|\psi(0)\rangle=|0\rangle^{\otimes N}$, and at $t=0$ we quench the transverse field through the quantum phase transition point from $\Omega=0$ to $\Omega=G$.
The main innovation of the present paper is to point out a surprising connection between DPTs in a quantum system and rainbows. Specifically, we show that in a paradigmatic spin model, the DPTs occur in a special window of times that is closely analogous to Alexander’s dark band (DB), which is the name given to the dark region in the sky between the primary and secondary bows of an ordinary optical rainbow, and is named after Alexander of Aphrodisias who described it around A.D. 200 [45]. In the context of the quantum spin system we call the analogoue of this phenomenon the quantum dark band. At the heart of our treatment is the observation that both Alexander’s DB and the quantum DB are bounded by caustics, which are singularities of the classical theory where the probability density diverges (in the rainbow case the two caustics are precisely the primary and secondary bows). Caustics have the special property of structural stability meaning that they are stable against perturbations and hence have universal qualitative properties that are independent of the quantitative details of the system under consideration [46, 47].
Our key finding is that the nonanalytic behaviour of the Loschmidt echo can be linked to a switching in exponential dominance of the evanescent tails from the two bounding caustics. This interpretation not only directly implies the structural stability of DPTs (via the structural stability of caustics) but can also be utilized for analytic calculations using methods such as Poisson resummation, and in this way we are able to obtain exact limiting expressions for the Loschmidt echo and to propose a measurement scheme for the DPTs with finite resources.
The universal properties of caustics can be explained by catastrophe theory [48, 49] (not to be confused with the orthogonality catastrophe, which is unrelated). This shows that structurally stable singularities are partitioned into distinct equivalence classes; in the two dimensions relevant to this work the two classes are folds and cusps, or $A_{2}$ and $A_{3}$ in the ADE classification scheme. Each equivalence class comes with its own set of scaling exponents under which the catastrophes display self-similar scaling. This is reminiscent of the universality classes and critical exponents of equilibrium phase transitions, except that in our treatment the non-analyticity occurs in the action rather than the free energy and is therefore intrinsically dynamical. Thus, the catastrophe theory approach naturally emphasizes the generic structure and universal aspects of the dynamics associated with DPTs. Furthermore, an appealing feature of catastrophe theory is that each class of catastrophe corresponds to a particular morphology (shape) of singularity and this allows for a particularly visual and hence intuitive interpretation of DPTs.
@@
The plan for the rest of this paper is as follows. In Section II we specify the model we use, which is the fully connected transverse field Ising model (FCTFIM), and Section III gives background information about the Fock state caustics that develop in the dynamics. Section IV gives an overview of our main results, including an explanation of how, in the vicinity of a DPT, the Loschmidt echo is analogous to the light intensity within Alexander’s DB, but with the role of time replacing the angle to the sky. Section V gives a more indepth treatment of rainbows, and Section VI presents the main theoretical calculation in the paper which gives the rate function in the semiclassical regime. Section VII discusses a possible measurement scheme, which is a tricky issue because the Loschmidt echo is exponentially small at the DPT. Our solution is to perform two different measurements on two halves of the system. We give a general discussion of switching lines which are the essence of DPTs in Section VIII, and our conclusions are set out in Section IX.
II The model
In this article we focus on the FCTFIM [44, 25] which is the fully-connected version of the well-known transverse-field Ising model (TFIM). This has been used as the minimal model for dynamics in a wide range of physical systems including bosonic Josephson junctions in ultracold atomic gases [19, 20, 21, 22, 23, 24], weak links in superfluid 3He [50], ferroelectrics such as KH2PO4 and ferromagnets such as Dy(C2H5SO4)39H2O [51], exciton polaritons in semiconductors [52, 53], and light beams carrying angular momentum [54]. The model is also relevant to TFIM systems that are not fully connected but have long range interactions such as those realized using trapped atomic ions [55, 56, 57, 38, 39], and the Dicke model describing cold atoms in optical cavities [58, 59, 60, 61]. Given that DPTs have been studied in the FCTFIM before [62, 44, 63, 64, 25], the goal of this paper is to establish a more intuitive explanation of the phenomenon in terms of caustics which are a generic feature of wave dynamics and which occur in this model in the real-time dynamics of the state in the Fock space representation.
@@ II The model
The FCTFIM [44, 25] consists of $N$ two-level systems (Pauli operators $\vec{\sigma}^{i}$) with Hamiltonian
@@
These are commonly known as the Dicke states. However, we will refer to them as Fock states because they are the symmetrized single particle eigenstates of the number difference operator (giving the difference between the number of up and down spins, i.e. the magnetization). For the sake of working with a specific example, in what follows we set $\Omega=G$, and choose the initial state to be a single Fock state $|\psi_{0}\rangle=|j,m_{0}\rangle$, with $m_{0}$ chosen as the closest integer (or half integer) to $0.6j$. Although a single Fock state might not appear to be a ‘typical’ initial condition, it is in fact not special because small variations in the initial state or the parameters in the Hamiltonian do not lead to qualitatively different outcomes in the wave function dynamics, as is to be expected from the structural stability of catastrophes. For example, choosing a narrow gaussian wavepacket of Fock states still leads to similar fold and cusp patterns [65, 66]. Likewise, going to the opposite situation of an initial phase state (which is very broad in Fock space) also gives rise to caustics [67], although of course they are in different positions. The key ingredient for caustics to appear is that the initial state must be energetically far above the ground state and such states can easily be generated in sudden quenches. Similar to the caustics, DPTs also occur generically in the FCTFIM for quenches from localized states (such as Fock states) to the regime with a strong transverse field (as for instance in Fig. 1).
@@
Figure 2: Left: The modulus of the exact quantum wave function plotted in Fock space as a function of time. The shading gives the magnitude of the overlap $|\langle j,m|\psi(t)\rangle|$ with each Fock state. The parameters are $\Omega=G$, $j=80$, and the initial state is the single Fock state $|j=80,m=48\rangle$. The envelopes separating the bright and dark regions are fold caustics, and these meet at cusp caustics (in order to faciliate comparison with optical rainbows we use the terminology ‘bright’ and ‘dark’ although in the image colourmap higher amplitudes correspond to blacker shades and lower amplitudes to whiter shades). In the full quantum theory shown here, folds are dressed by Airy functions whose characteristic fringe pattern can be seen near the edges of the shaded (‘bright’) regions. Cusps are dressed by Pearcey functions as occurs near $m/j=-0.6$ and time $\Omega t\approx 2.4$. Right: The Loschmidt echo $L(t)$, or return probability, is given by the cut through the general wave function corresponding to the initial state. The cut is depicted by the horizontal blue line in the left panel. The Loschmidt echo shown here features a one dimensional slice through two fold caustics that bound a dark forbidden region where the wave function is exponentially small.
III Caustics in Fock space
Caustics are usually associated with waves or rays in real space and are the most intense regions of a wave field. In geometric optics they are the envelopes of families of rays and form quite generically through natural focusing, e.g. optical rainbows focused by water droplets in the sky and gravitational lensing by mass distributions in the cosmos [46]. Similarly, freak waves, tidal bores and tsunamis are caustics resulting from the natural focusing of water waves by the topography of the sea bed or estuary shape [68, 69, 70, 71, 72]. Caustics are singularities of the classical theory: they are regions where the wave amplitude diverges in the short wavelength (ray) limit $\lambda\rightarrow 0$, and their natural mathematical description is in terms of catastrophe theory [46, 47]. Caustics in Fock space are the extension of these ideas to many-body quantum mechanics. A single ray in Fock space represents a single classical field configuration, while quantum configurations can be represented by the superposition of many rays. Whereas ordinary caustics are resolved by going to the wave theory where interference removes the singularity, caustics in Fock space are resolved by second-quantization where the discreteness of field excitations is the new feature needed to remove the singularity. The theory of Fock-space caustics has been developed over the past decade in the context of sudden quenches in cold atom systems [73, 74, 65, 75, 67] and is also relevant to dissipative systems [76].
Catastrophe theory makes the remarkable prediction that there are only a finite number of distinct shapes that structurally stable singularities can take, depending on the dimension of the space in which they live [48, 49]. In the two dimensions which are relevant to the dynamics of the FCTFIM, namely the one dimension of Fock space that gives the total magnetization (difference between the number of up and down spins) plus time, the only structurally stable catastrophes are fold lines and cusp points. Furthermore, in the wave theory each type of catastrophe is dressed by its own characteristic local wave function known as a ‘diffraction integral’ or ‘wave catastrophe’ [77, 78]. For the fold this is the Airy function and for the cusp it is the Pearcey function [79]. Both of these wave functions can be seen occurring in the left hand panel of Fig. 2 that shows the development of a many-body wave function in Fock space as a function of time. Higher dimensional models display higher wave catastrophes [67] that form a hierarchy with the higher ones containing the lower ones. Each type of catastrophe forms an equivalence class with its own scaling properties governed by exponents known as Arnold and Berry indices [47].
In fully-connected spin models the dynamics in Fock space are specified by the set of quantum amplitudes
$\{\langle j,-j|\psi(t)\rangle,\langle j,-j+1|\psi(t)\rangle,\ldots,\langle j,j|\psi(t)\rangle\}$ (6)